The President of UTA is interested in estimating the proportion of Statistics students who drop BUSA 3321 this semester with an UPPER ONE-TAILED 100*(1-α)% confidence interval (α is shown on right). This semester there were n students enrolled in 3321 (shown on right) and Xo of those students dropped (shown on right). Calculate the UPPER ONE-TAILED confidence limit for the true proportion who drop 3321. Round all Z scores off to 2 decimals.

α = 0.05
n = 110
Xo = 42
b 0.458
d 0.473
c 0.441
a 0.306
e None
Answer #1

sample proportion, = 0.3818
sample size, n = 110
Standard error, SE = sqrt(pcap * (1 - pcap)/n)
SE = sqrt(0.3818 * (1 - 0.3818)/110) = 0.0463

Given CI level is 90%, hence α = 1 - 0.9 = 0.1
α/2 = 0.1/2 = 0.05, Zc = Z(α/2) = 1.64

upper one tailed CI = pcap + z*SE
= 0.3818 + 1.64 * 0.0463
= 0.458